# references.bib A Learning-based Quadcopter Controller with Extreme Adaptation

### v1

@@ VI Conclusion
This work demonstrates how a single adaptive controller can effectively bridge the gap between high-level planning and the intricate physical dynamics by adapting to model disparities between quadcopters down to the motor level. Our design focuses on creating a low-level controller intended to replace traditional low-level quadcopter controllers, thereby eliminating the need for accurate model estimation and iterative parameter tuning. Our approach leverages a combination of imitation learning from model-based controllers and reinforcement learning to address the challenges of training a sensor-to-actuator controller at high frequencies. The introduction of an instant reward feedback ensures that the controller remains responsive and agile. In addition, we develop a quadcopter randomization method during training that aligns with real-world constraints, further enhancing its adaptability. The controller’s ability to estimate a latent representation of system parameters from sensor-action history, along with realistic domain randomization, empowers it to generalize across a broad spectrum of quadcopter dynamics. This capability extends to unseen parameters, with an adaptation range up to 16 times broader than the training set. The single policy trained solely in simulation can be deployed zero-shot to real-world quadcopters with vastly different designs and hardware characteristics. It also demonstrates rapid adaptation to unknown disturbances, such as off-center payloads, wind, and partial motor failure. These results highlight the potential of our approach for extreme adaptation for drones and other robotic systems, while enabling robust control in the face of real-world uncertainties.
VII Acknowledgement
@@
This work was supported by the Hong Kong Center for Logistics Robotics, the DARPA Transfer from Imprecise and Abstract Models to Autonomous Technologies (TIAMAT) program, the Graduate Division Block Grant of the Dept. of Mechanical Engineering, UC Berkeley, and ONR MURI award N00014-21-1-2801. The authors would like to thank Ruiqi Zhang and Teaya Yang for their help with the experiments. The experimental testbed at the HiPeRLab is the result of contributions of many people, a full list of which can be found at hiperlab.berkeley.edu/members/.

### v2

@@ VII Conclusion
This work demonstrates how a single adaptive controller can effectively bridge the gap between high-level planning and the intricate physical dynamics by adapting to model disparities between quadcopters down to the motor level. Our design focuses on creating a low-level controller intended to replace traditional low-level quadcopter controllers, thereby eliminating the need for accurate model estimation and iterative parameter tuning. Our approach leverages a combination of imitation learning from model-based controllers and reinforcement learning to address the challenges of training a sensor-to-actuator controller at high frequencies. The introduction of an instant reward feedback ensures that the controller remains responsive and agile. In addition, we develop a quadcopter randomization method during training that aligns with real-world constraints, further enhancing its adaptability. The controller’s ability to estimate a latent representation of system parameters from sensor-action history, along with realistic domain randomization, empowers it to generalize across a broad spectrum of quadcopter dynamics. This capability extends to unseen parameters, with an adaptation range up to 16 times broader than the training set. The single policy trained solely in simulation can be deployed zero-shot to real-world quadcopters with vastly different designs and hardware characteristics. It also demonstrates rapid adaptation to unknown disturbances, such as off-center payloads, wind, and loss of efficiency in motors. These results highlight the potential of our approach for extreme adaptation for drones and other robotic systems, while enabling robust control in the face of real-world uncertainties.
VIII Acknowledgement
@@
This work was supported by the Hong Kong Center for Logistics Robotics, the DARPA Transfer from Imprecise and Abstract Models to Autonomous Technologies (TIAMAT) program, the Graduate Division Block Grant of the Dept. of Mechanical Engineering, UC Berkeley, and ONR MURI award N00014-21-1-2801. The authors would like to thank Ruiqi Zhang and Teaya Yang for their help with the experiments. The authors also would like to thank Daniel Gehrig for his valuable input in the ablation analysis of our method. The experimental testbed at the HiPeRLab is the result of contributions of many people, a full list of which can be found at hiperlab.berkeley.edu/members/.
Appendix A Baseline Implementation
A detailed explanation of each baseline’s implementation is provided to support reproducibility. The baselines are structured into high-level and low-level controllers. Most baseline names follow the format high-level controller-low-level controller, with exceptions explicitly noted. For instance, PID-PD∗ indicates that PID serves as the high-level controller, while PD∗ represents the low-level controller. Identical control names imply the same implementation, and explanations are omitted if they were already described for previous baselines.
A-A PID-PD∗
A-A1 PID
The high-level PID controller is implemented as described in Section III. It generates high-level commands, including body rates and mass-normalized collective thrust, for other low-level controllers to track.
A-A2 PD∗
This baseline matches the expert controller used during training, as outlined in Section II-D. Unlike other baselines, it has access to ground-truth model parameters for control calculations. Consequently, it serves as both an expert reference and a performance upper bound, providing context for evaluation.
A-B PID-PDn
A-B1 PDn
This baseline is similar to PD∗ but calculates control using only the nominal model . As such, it represents a performance lower bound for comparison.
A-C PID-
A-C1 at Low Level
We implement the adaptive controller as an augmentation to the PID-PDn control, applying the adaptive controller at the low level. Beyond serving as a baseline, it also acts as a counterpart to -PDn, which applies the same adaptation within the high-level control structure. The comparison between the two methods can validate our design choice of control hierarchy. The implementation is same to [10] and the controller’s block diagram is shown in Figure 13. The control receives the desired motor thrust from PDn and computes the augmented thrust using the control and adaptation law defined in (6-13) in [10]. The resulting thrust command is then converted into motor speed. Unless otherwise specified, all model constants used for control calculations in this framework and in all baselines, except PD∗, are derived from the nominal model.
A-D -PDn
A-D1 at High Level
We implement a similar controller as described previously but the uncertainty is compensated at the level of thrust and body torque. We refer to the framework in [47] in which the controller gives thrust and torque command. The control and adaptation law is defined in (19-24) in [47]. Figure 14 outlines the block diagram of the controller.
A-E Geo-A
This baseline is a geometric adaptive controller both on the high level and low level, abbreviated as Geo-A, an exception to our naming convention. We choose Geo-A as a baseline to contextualize our proposed approach. This method ensures precise trajectory tracking and robust stabilization for quadrotors by adapting to uncertainties and disturbances in a geometrically consistent manner.
We implement the geometric adaptive controller based on [12], replacing the original geometric attitude controller with a tilt-prioritized control from [48] to enable tracking of dynamically infeasible trajectories. Figure 15 provides an overview. The position controller, based on (23-25) in [12], computes target acceleration, which is decomposed into target thrust and target attitude using (14-17) in [48]. The attitude is processed by tilt-prioritized control ((25-28) in [48]), and the resulting angular velocity and thrust are sent to the angular velocity controller ((15-16) in [12]). Finally, thrust commands are mapped to motor speeds.
A-F PID-INDI-A
A-F1 INDI-A
We implement the adaptive INDI controller based on [13] and the block diagram is detailed in Figure 16. INDI is a sensor-based control which uses instant sensor measurement to represent the system dynamics so that to be more robust to unmodelled uncertainties in the rotation dynamics. To enhance its adaptiveness, [13] introduces onboard adaptive parameter estimation to update control effectiveness.
The original implementation requires measurements of angular acceleration, motor speed, and motor speed acceleration. The setup provides more informative observations than other baselines and our method, which do not have access to motor feedback. For a fair comparison, we modify the control law to reduce reliance on motor feedback, using only motor speed estimates obtained through a first-order response model with a time constant of 10ms while preserving the original linear update adaptation law.
After knowing the desired mass-normalized thrust and angular velocity from the high-level controller, we can compute the desired torque using (6). Then we can compute the desired motor speed command using (7) and (8), yielding:
We want to use instantaneous sensor feedback to compute the desired motor speed to be against model uncertainty and disturbance. Therefore, we decompose the thrust and torque into current measurement and desired value (see (30) and (31) of [49] for detailed derivations of torque decomposition).
So that the effect of unmodeled rotational dynamics is captured by the measurement of angular acceleration and torque . Note the desired angular acceleration is gained through (5).
Similarly, the desired motor speed is decomposed as:
By rearranging the terms, we obtain the following equation to solve for :
This equation can then be reformulated into the INDI framework by introducing the control allocation matrices and :
with the adaptation law defined as
where denotes the difference between the current and previous sampled variables. A more detailed analysis on the stability and performance of this method could be found in [13].